Understanding the optimal capital structure Review this situation: Universal Exports Inc. is trying to identify its optimal capital structure. Universal Exports Inc. has gathered the following financial information to help with the analysis. Debt Ratio Equity Ratio EPS DPS Stock Price 30% 40% 50% 60% 70% 7096 1.25 0.55 36.25 6096 1.40 0.60 37.75 50% 1.60 0.65 39.50 40% 1.85 0.75 38.75 30% 1.75 0.70 38.25 Which capital structure shown in the preceding table is Universal Exports Inc. s optimal capital structure? O Debt ratio-40%; equity ratio-6096 O Debt ratio-60%; equity ratio-40% O Debt ratio-30%; equity ratio-70% @) Debt ratio-50%; equity ratio-50% O Debt ratio-70%; equity ratio-30% Consider this case: Globex Corp. is an all-equity firm, and it has a beta of 1 . It is considering changing its capital structure to 60% equity and 40% debt. The firms cost of debt will be 10%, and it will face a tax rate of 35%. What will Globex Corp.s beta be if it decides to make this change in its capital structure?Now consider the case of another company: U.S. Robotics Inc. has a current capital structure of 30% debt and 70% equity. Its current before-tax cost of debt is 1096, and its tax rate is 35%. It currently has a levered beta of 1.25. The risk-free rate is 3%, and the risk premium on the market is 7.5% U.S. Robotics Inc. is considering changing its capital structure to 60% debt and 40% equity. Increasing the firms level of debt will cause its before-tax cost of debt to increase to 12%. Use the Hamada equation to unlever and relever the beta for the new level of debt. What will the firms weighted average cost of capital (WACC) be if it makes this change in its capital structure? Which of the following statements regarding a firms optimal capital structure are true? Check all that apply. The optimal capital structure minimizes the firms cost of debt. The optimal capital structure minimizes the firms WACC. The optimal capital structure maximizes the firms EPS. The optimal capital structure minimizes the firms cost of equity. The optimal capital structure maximizes the firms stock price
Answer #1

1.

Debt ratio=50%, Equity ratio=50%

2.

=1*(1+(1-35%)*40%/60%)=1.433

3.

=12%*(1-35%)*60%+40%*(3%+7.5%*1.25/(1+(1-35%)*30%/70%)*(1+(1-35%)*60%/40%))=11.6725%

4.

Option B,E